The region R is bounded by the curves y = 22 +1 and y = 3x - 1. Set up an integral blume of the solid obtained by rotating R calculate volume
Answer #1

we are given equations as

y=x^2+1

y=3x-1

It is rotated about x=4 line

Firstly, we can draw graph

Bound is from 1 to 2

we can use shell method

V = 2arhda

h=(3x-1)-(x^2+1)

r=4-x

now, we can set up integral

V=\int_{1}^{2}2\pi (4-x)((3x-1)-(x^2+1))dx

=2\pi \cdot \int _1^2x^3-7x^2+14x-8dx

=2\pi \left(\int _1^2x^3dx-\int _1^27x^2dx+\int _1^214xdx-\int _1^28dx\right)

we can solve each integrals

and then combine them

=2\pi \left(\frac{15}{4}-\frac{49}{3}+21-8\right)

V=\frac{5\pi }{6}............Answer